7.Police suspect that a van was breaking the speed limit when it collided with a parked car. Investigations provided them with the following data: mass of parked car                  = 800 kg,  mass of van = 1000 kg,  velocity of  car after collision = 5 m/s,  velocity of van after collision = 11 m/s


Show whether the van was breaking the 30 mph speed limit (equivalent to 13 m/s).


8.A dodgem car of mass 200 kg moves with a velocity of 3 m/s. The moving dodgem collides with a stationary one of mass 180 kg.


After the collision, the 200 kg dodgem continues to move in the same direction at a speed of 1 m/s.


Calculate the speed v with which the 180 kg (second) dodgem moves to the right after the collision.





1
Expert's answer
2021-05-16T17:59:04-0400

"7."

"\text{van: }"

"m_1 =1000 ;\vec u_1-?;\vec v_1= 11"

"\text{car:}"

"m_2 =800;\vec u_2=0;\vec v_2= 5"

"\vec u_1,\vec u_2\ \text{pre-collision speed}"

"\vec v_1, \vec v_2\ \text{speed after collision}"

"\text{according to the law of conservation of momentum:}"

"\vec p= m_1\vec u_1+m_2\vec u_2\ = m_1\vec u_1=1000\vec u_1"

"\vec p= m_1\vec v_1+m_2\vec v_2\ = 1000*11+800*5=15000"

"1000\vec u_1= 15000"

"\vec u_1= 15\frac{m}{s}>13\frac{m}{s}"

"\text{Answer: The van driver has exceeded the speed limit}"


"8."

"\text{dodgem car first:}"

"m_1 =200 ;\vec u_1=3;\vec v_1= 1"

"\text{dodgem car second:}"

"m_2 =180 ;\vec u_2=0;\vec v_2-?"

"\vec u_1,\vec u_2\ \text{pre-collision speed}"

"\vec v_1, \vec v_2\ \text{speed after collision}"

"\text{according to the law of conservation of momentum:}"

"\vec p= m_1\vec u_1+m_2\vec u_2\ = 200*3=600"

"\vec p= m_1\vec v_1+m_2\vec v_2= 200*1+180\vec v_2"

"200*1+180\vec v_2 =600"

"\vec v_2=\frac{400}{180}\approx2.22\frac{m}{s}"

"\text{Answer: }\approx2.22\frac{m}{s}"